OpenStudy (anonymous):

Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find b if a = 21 and c = 29.

OpenStudy (raden):

b^2 = c^2 - a^2

OpenStudy (anonymous):

Answers ; 40, 10 , 100, 20 ??

OpenStudy (raden):

b^2 = c^2 - a^2 b^2 = 29^2 - 21^2 solve for b

OpenStudy (anonymous):

ok but what is ^ ?

OpenStudy (raden):

^ is the exponent sign

OpenStudy (anonymous):

ohhhh.! okay gracias

OpenStudy (raden):

did you get it ?

OpenStudy (anonymous):

yes its 400

OpenStudy (raden):

that's just the value of b^2 b^2 = 400 b = .... ?

OpenStudy (anonymous):

ok now i just got confused,

OpenStudy (raden):